OpenStudy (anonymous):

A group of students and teachers took the bus to the Orange County Museum.Adult single-day passes cost $4.50 and student single-day passes cost $2.25. The total cost of all single-day passes was $117. How many adult passes were purchased?

OpenStudy (anonymous):

It doesnt say how many people?

OpenStudy (anonymous):

We don't have enough information right now to say.

OpenStudy (anonymous):

sorry it is 45 students

OpenStudy (anonymous):

Ok, we have two equations:\[S+T=45\]\[2.25S+4.5T=117\](I'm using S for the number of students and T for the number of teachers). Let's use elimination to solve for one of the variables. This means we're going to take one equation, multiply it by some number, and then subtract it from the other equation. Let's subtract 4.5 times the first equation from the second equation:\[2.25S+4.5T-4.5(S+T)=117-4.5 \times 45\]\[2.25S+4.5T-4.5S-4.5T=117-202.5\]\[-2.25S=-85.5\]\[\frac{-2.25S}{-2.25}=\frac{-85.5}{-2.25}\]\[S=38\]So that tells us that 38 students were on the trip. You can now take that value for S to get the number of teachers. :)

OpenStudy (anonymous):

But it would probably be better practice for you if you did what I did, except trying to eliminate S this time.

OpenStudy (anonymous):

i got 7 is that right

OpenStudy (anonymous):

@Aylin

OpenStudy (anonymous):

Yes that's right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Luis_Rivera Your methodology is flawed. First, you have no way of knowing that the ratio between students and teachers is 1/2. You only know that a student ticket costs 1/2 as much as a teacher ticket. Additionally, you then apparently claim that the number of students is 13, yet this information was not given. Plus, 13+26 is NOT 45, it is 39. Your solution fails to take into account the total number of students and teachers. It would behoove you to understand the problem before you attempt to "help".

OpenStudy (anonymous):

um that is not an option

OpenStudy (anonymous):

n both of u thank u for ur help n sorry to be rude but go somewhere else with that cuz i dont got time to be flipping back in forth to see ppl bickering like children i asked his question for help not to see ppl fight over the answer u wanna tell each other its wrong then do it like an adult n talk it over like dang i did not sign up for this to babysit so guys chill out n work out yall stinking differences Thank You Very Much LOVE ME!!!!!!!!!!!! MWAH THANK U BOTH FOR UR HELP BUT PLEASE DONT BICKER ON MY QUESTION'S ANYMORE THANK YOU I LOOK FORWARD TO HELPING U GYS N WORKING WITH U GUYS N THE FUTURE!!!!!!!!!